# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset (GIST)

### v4

@@ 6.2 Experiment Setup
We conduct experiments on two evaluation sets: the 60-60 Evaluation Set and the AI Papers and Model Cards Evaluation Set. The first evaluation set is from the 60-60 initiative, providing terminology translations into five target languages. Based on our analysis in Section 4.2, which indicates that our dataset offers higher quality than the 60-60 set, we enhance their labels by replacing AI terminology in their translations with GPT-4o while maintaining grammatical correctness and ensuring no loss in translation quality. This process generates updated ground truth labels for the evaluation.
The second evaluation set is manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by (Liu et al., 2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we use Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompt GPT-4o to refine these translations by updating AI terminology. This evaluation set is designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards.
@@
We evaluate the following models: gpt-4o-mini (Hurst et al., 2024), hf-seamless-m4t-large (Barrault et al., 2023), nllb-200-3.3B (Costa-jussà et al., 2022), aya-23-8B (Aryabumi et al., 2024), and aya-expanse-8B (Dang et al., 2024). While the prompting method applies to all models, the word alignment and decoding methods are not applicable to gpt-4o-mini, as it does not provide access to its model weights.
@@
We adopt BLEU (Papineni et al., 2002) and COMET metrics (Rei et al., 2020) for the quantitative evaluation of translation quality, following the methodology of Salesky et al. (2023). We utilize the wmt22-comet-da model (Rei et al., 2022).
@@ 6.3 Experiment Results
Figure 3: The interface demonstrates a paper example from (Tang et al., 2024) and introduces a new translation feature on the ACL Anthology website. Non-English speakers can click the “Translate” button at the bottom right of the webpage and select their preferred language for translation. The page dynamically displays two translations side by side: one using direct translation from a machine translation model (e.g., seamless in the figure) and the other enhanced with prompting-powered refinement applied to the left translation. Highlighted text indicates updated terminology integrated from GIST, showcasing improved translation performance.
@@
Several trends emerge when comparing different terminology integration approaches. The prompting-powered refinement method consistently outperforms direct translation across nearly all languages, models, and evaluation metrics, which highlights its effectiveness in effectively incorporating AI-specific terminology into translations.
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese exhibits minimal morphological changes, allowing straightforward substitution of terminology with limited disruption to surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more complex and error-prone. By comparing the fourth and second values in each column of Table 2, we observe that the post-hoc prompting ablation enhances translation scores, aligning them with the performance of the prompting-powered method across all languages and models, thereby supporting our claims.
We also conduct a qualitative manual evaluation of the generation results produced by constrained beam search and logit adjustment methods. Both approaches exhibit extremely slow performance, running approximately 100 times slower than post-hoc methods. Furthermore, forcing specific word ids to appear in the output did not perform well, suffering from similar issues as the word alignment method in disrupting syntactical dependencies and morphological agreements. During manual inspection, we observe that constrained beam search behaves similarly to the logit adjustment method with a large scaling factor: the generated sentences often failed to maintain proper grammar, and terminological terms were frequently repeated multiple times within the same output. Conversely, with a small scaling factor in the logit adjustment method, the terms were often omitted entirely, demonstrating little to no effect on the output. These findings highlight the limitations in these traditional decoding approaches compared to the prompt-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages.
@@ 7 Website Demonstration
To facilitate real-world usage of our AI terminology dictionary, we built on the 60-60 initiative by modifying the ACL Anthology website layout and introducing a new terminology translation feature, [^6] https://acl6060.org/ as illustrated in Figure 3. The ACL Anthology website was chosen for this demonstration due to its extensive collection of AI-related research papers, making it an ideal platform to showcase the potential impact of our work. In this demonstration, translations of key terms are refined and standardized based on the terminology dictionary we developed, providing more accurate terminology translations.777 Our code for the website is available at https://github.com/ImanOu123/acl-anthology-mod. This enhancement represents a step toward improving access to AI knowledge for non-English speakers by offering a consistent and reliable translation system, and broadens the accessibility of AI research to a global audience.
@@ 8 Conclusion
We present GIST, a large-scale multilingual AI terminology dataset addressing gaps in translating AI-specific terms. Combining LLM-based extraction and validation with human expertise, it includes 5K English terms with translations into Arabic, Chinese, French, Japanese, and Russian, surpassing the ACL 60-60 benchmark. LLM prompting proved effective for post-hoc terminology integration, improving translation quality across metrics like BLEU and COMET. We also provide a website demonstration to enhance the accessibility of our work for non-English speakers, supporting equitable AI knowledge access and fostering global collaboration in AI research.
@@
The evaluation set was created using Google Translate and GPT-4, tools whose translation accuracy has not been verified in our work by human experts. This reliance might introduce potential errors in the evaluation data, affecting its reliability.
@@
In crowdsourced experiments, we respected participant privacy by not collecting any demographic information. All contributors were fairly compensated, with the total cost for annotations amounting to approximately \$1,000. Additionally, our collected terminology dataset does not involve ethically sensitive or controversial content, focusing exclusively on technical terms relevant to AI.
@@ Acknowledgments
We extend our gratitude to the Minister of State for Artificial Intelligence of the United Arab Emirates for making the Arabic AI terminology dictionary publicly available. We thank DataFranca for granting us access to their French AI terminology dictionary. We thank Jiqizhixin for sharing their Chinese AI terminology database. We also thank the Russian Association of Artificial Intelligence for providing access to the Russian AI terminology dictionary. For more details, please refer to Section A.2.
Appendix A Additional Dataset Details
A.1 Additional Dataset Statistics
@@
Conclusion.
The statistical analysis highlights the diversity and interdisciplinary nature of the dataset. Figures 8 and 9 illustrate domain-wise distributions and semantic clusters, while Table 3 quantifies lexical variations across languages. Together, these findings provide a robust understanding of the dataset’s structure, supporting its utility for multilingual and domain-specific AI applications.
A.2 Terminology Collection Details
We integrate terms from the WikiPedia Glossary of AI888en.wikipedia.org/wiki/Glossary_of_artificial_intelligence, which serves as a comprehensive starting point for artificial intelligence (AI)-related terminology, ensuring alignment with globally recognized concepts and definitions. To enhance multilingual coverage and domain relevance, we also include terms from several other specialized AI terminology initiatives across different languages:
Arabic AI Dictionary999ai.gov.ae/ar/ai-dictionary/: Published by the Arabic Government AI Office, this resource aims to elevate the status of the Arabic language in AI, standardize terminology, reduce linguistic ambiguity, and foster better integration of Arabic speakers into the global AI community. By providing accurate translations and clear definitions for English AI terms, this dictionary promotes knowledge dissemination and encourages collaboration within the Arabic-speaking AI ecosystem.
GRAND_LEXIQUE_FRAN%C3%87AIS
Russian AI Dictionary [^12] https://www.raai.org/pages/UGFnZVR5cGU6MTAwMw==: This initiative captures the interdisciplinary nature of AI by including terminology drawn from fields such as logic, psychology, linguistics, and cybernetics. Leveraging contributions from Russian and Soviet experts, this dictionary emphasizes the frequency and relevance of terms within AI-specific contexts, providing a culturally adapted yet globally aligned resource for Russian-speaking researchers.
A.3 Terminology Translation Details
Terminology Dataset Collection
Category Selected Conferences
AI AAAI, IJCAI
CV CVPR, ECCV, ICCV
ML ICLR, ICML, NeurIPS, KDD
NLP ACL, EMNLP, NAACL, EACL
LREC, COLING, CoNLL
Web & IR SIGIR, WWW
Table 4: Top AI conferences included in our dataset collection.
Prompt Design for Translation
Arabic Chinese French Japanese Russian
Three-Model Agreement Ratio 10.11% 42.71% 9.86% 16.60% 17.93%
Two-Model Agreement Ratio 30.52% 36.82% 45.15% 40.44% 38.36%
Translation Agreement Analysis
We evaluated the consistency of translations across three models: Claude 3 Sonnet, GPT-3.5-Turbo, and Google Translate API. Table 5 summarizes the agreement ratios for five target languages: Arabic, Chinese, French, Japanese, and Russian. The results indicate significant variation in agreement ratios across languages, with Chinese achieving the highest three-model agreement (42.71%) and Arabic having the lowest (10.11%). The detailed discussion is in Section 3.2.
Human Translation via Mturk
To further validate the translations, we employed human translator through the MTurk platform. Each terminology was translated by 10 independent annotators, resulting in a set of 10 translation candidates for each term. The platform layout and guidance for helping translators give these translations is shown in Figure 6.
A.4 Dataset Quality Assessment Details
To evaluate the quality of translations in our dataset, we conducted two tasks involving human annotators. Annotators were presented with pairs of translations generated by different methods and were tasked with evaluating their relative quality. The analysis of their ratings is summarized in Tables 6, 7, and 8. Below, we describe the findings in detail.
In Task 1, annotators compared translations selected by GPT-4o (Method 1) with majority-voted translations generated by human annotators (Method 2). The choices included:
Arabic 739, 45.76% 461, 28.54% 329, 20.37% 72, 4.46%
Chinese 468, 50.59% 266, 28.76% 163, 17.62% 25, 2.70%
French 438, 48.67% 274, 30.44% 170, 18.89% 17, 1.89%
Japanese 587, 56.99% 251, 24.37% 159, 15.44% 27, 2.62%
Russian 626, 54.43% 348, 30.26% 150, 13.04% 14, 1.22%
Arabic 376, 46.42% 238, 29.38% 143, 17.65% 46, 5.68%
Chinese 197, 37.17% 228, 43.02% 85, 16.04% 17, 3.21%
French 152, 39.48% 168, 43.64% 53, 13.77% 10, 2.60%
Japanese 295, 57.28% 162, 31.46% 36, 6.99% 21, 4.08%
Russian 172, 39.09% 198, 45.00% 38, 8.64% 25, 5.68%
In Task 2, annotators compared translations from our dataset (Method 1) with those in the 60-60 initiative evaluation set (Method 2). The distribution of choices is shown in Table 7. In this task, a higher percentage of annotators preferred translations from Method 2 (category B) for most languages, especially in Chinese (43.02%), French (43.64%), and Russian (45.00%). However, Japanese translations from Method 1 had a significantly higher percentage in category A, with 57.28% of annotators agreeing that both translations were good. Instances where both translations were rated as bad (category D) remained low across all languages, ranging from 2.60% to 5.68%.
Arabic Chinese French Japanese Russian
Task 1 0.30 0.22 0.37 0.21 0.39
Task 2 0.22 0.41 0.50 0.39 0.41
Inter-Annotator Agreement
Conclusion
A.5 Dataset Coverage Assessment Details
Domain Distribution.
Semantic Clustering.
Coverage Analysis through Rarefaction.
Figure 8: Terminology distribution of the top-6 AI domains in GIST.
Figure 9: UMap visualization of terms in GIST by domain.
A.6 Terminology Examples
@@ 50 K-means 100 Precision 150 Coreference resolution
Appendix B Additional Experiment Details
B.1 Terminology Refinement via Prompting Details

### v5

@@ 6.2 Experiment Setup
We conduct experiments on two evaluation sets: the 60-60 Evaluation Set and the AI Papers and Model Cards Evaluation Set. The first evaluation set is from the 60-60 initiative, providing terminology translations into five target languages. As detailed in Section 4.2, our analysis shows that our dataset achieves higher quality compared to the 60-60 set. Motivated by this, we enhance the labels of the 60-60 evaluation set by replacing AI terminology in their translations with GPT-4o while maintaining grammatical correctness and ensuring no loss in translation quality. This process generates updated ground truth labels for the evaluation.
The second evaluation set is manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by Liu et al. (2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we use Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompt GPT-4o to refine these translations by updating AI terminology. This evaluation set is designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards. Both evaluation set references have been verified by human experts and exhibit high quality. Further details are provided in Section C.4.
@@
We evaluate the following models: gpt-4o-mini (Hurst et al., 2024), hf-seamless-m4t-large (Barrault et al., 2023), nllb-200-3.3B (Costa-jussà et al., 2022), aya-23-8B (Aryabumi et al., 2024), and aya-expanse-8B (Dang et al., 2024). While the prompting method applies to all models, the word alignment method is not applicable to gpt-4o-mini, as it does not provide access to its model weights.
@@
We adopt BLEU (Papineni et al., 2002), COMET (Rei et al., 2020), ChrF (Popović, 2015), ChrF++ (Popović, 2017), and TER metrics (Snover et al., 2006) for the quantitative evaluation of translation quality, following the methodology of Salesky et al. (2023). We utilize the wmt22-comet-da model (Rei et al., 2022) to compute COMET scores.
@@ 6.3 Experiment Results
Figure 3: The interface demonstrates a paper example from Tang et al. (2024) and introduces a new translation feature on the ACL Anthology website. Non-English speakers can click the “Translate” button at the bottom right of the webpage and select their preferred language for translation. The page dynamically displays two translations side by side: one using direct translation from a machine translation model (e.g., seamless in the figure) and the other enhanced with prompting-powered refinement applied to the left translation. Highlighted text indicates updated terminology integrated from GIST, showcasing improved translation performance.
@@
Several trends emerge when comparing different terminology integration approaches. The prompting-powered refinement method consistently outperforms direct translation across nearly all languages, models, and evaluation metrics, which highlights its effectiveness in incorporating AI-specific terminology into translations.
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese exhibits minimal morphological changes, allowing straightforward substitution of terminology with limited disruption to surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more complex and error-prone. These findings underscore the superior performance of the prompt-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages. See Sections D.1 and D.2 for further discussion of the results. We further validate our findings using a one-sided paired t-test in Section D.3.
@@ 7 Website Demonstration
To facilitate real-world usage of our AI terminology dictionary, we built on the 60-60 initiative by modifying the ACL Anthology website layout and introducing a new terminology translation feature, [^6] https://acl6060.org/ as illustrated in Figure 3. The ACL Anthology website was chosen for this demonstration due to its extensive collection of AI-related research papers, making it an ideal platform to showcase the potential impact of our work. In this demonstration, translations of key terms are refined and standardized based on the terminology dictionary we developed, providing more accurate terminology translations.This enhancement represents a step toward improving access to AI knowledge for non-English speakers by offering a consistent and reliable translation system, and broadens the accessibility of AI research to a global audience.
@@ 8 Conclusion
We present GIST, a large-scale multilingual AI terminology dataset addressing gaps in translating AI-specific terms. Combining LLM-based extraction and validation with human expertise, it includes 5K English terms with translations into five languages, surpassing the ACL 60-60 benchmark. LLM prompting proved effective for post-hoc terminology integration, improving translation quality across five metrics. We also provide a website demonstration to enhance the accessibility of our work for non-English speakers, supporting equitable AI knowledge access and fostering global collaboration in AI research.
@@
Furthermore, while our methodology is tailored for AI terminology translation, its application to other domains may require adaptation. Although our LLM + Human hybrid framework for data collection, translation, and evaluation is broadly applicable and does not rely on AI-specific models or tools, domain-specific terminology translation poses unique challenges, such as variations in terminology collection and the need for domain-specific expertise in evaluation.
@@
In crowdsourced experiments, we respected participant privacy by not collecting any demographic information. All contributors were fairly compensated according to MTurk’s payment standards. Additionally, our collected terminology dataset does not involve ethically sensitive or controversial content, focusing exclusively on technical terms relevant to AI.
@@ Acknowledgments
We extend our gratitude to the Minister of State for Artificial Intelligence of the United Arab Emirates for making the Arabic AI terminology dictionary publicly available. We thank DataFranca for granting us access to their French AI terminology dictionary. We thank Jiqizhixin for sharing their Chinese AI terminology database. We also thank the Russian Association of Artificial Intelligence for providing access to the Russian AI terminology dictionary. For more details, please refer to Section B.2.
Appendix A Further Discussion on AI Terminology Translation
In many non-English texts nowadays, AI terminology is often borrowed as untranslated loanwords. However, translating these terms into local languages is crucial for the following reasons:
  1. 1.
    Translation democratizes AI for non-English speakers. Translating AI terminology into native languages plays a vital role in making the field more accessible to non-English-speaking communities. For those outside the AI field, translating these terms fosters a clearer understanding, as it is easier for them to grasp complex concepts in their own language rather than through borrowed English terms. Prior studies highlight that translating English loanwords into Arabic and other languages helps to fill lexical gaps (Calabrese and Wetzels, 2009; Al-Athwary and Ali, 2024), making terms more meaningful while preserving linguistic and cultural identities. This practice promotes the standardization of AI terminology, reduces dependency on foreign languages, and improves educational accessibility for non-English speakers.
  2. 2.
    Our work aligns with and extends the ACL 60-60 initiative (Salesky et al., 2023), which aims to translate AI-specific terms to enhance multilingual scientific communication and foster AI inclusivity. By expanding this effort, we seek to improve the scale, scope, and quality of AI terminology translation, furthering the global reach of AI research.
  3. 3.
    The prevalence of untranslated loanwords varies significantly across languages. The prevalence of untranslated loanwords is more common in languages influenced by English or with linguistic similarities to English. However, this is not the case for many other languages. As a concession to this variability, we employed GPT-4o-mini to determine whether terms in the GIST dataset should be translated. Table 3 presents the results, based on the prompt in Figure 4, which highlights the disparities across languages.
Language Percentage of Terms Requiring Translation
Chinese 95.0%
Arabic 73.0%
French 72.0%
Japanese 53.8%
Russian 80.5%
Table 3: Proportion of AI terms requiring translation across different languages.
Figure 4: Prompt used to determine whether GIST terms should be translated.
Appendix B Additional Dataset Details
B.1 Additional Dataset Statistics
@@
Terminology Examples
@@ 50 K-means 100 Precision 150 Coreference resolution
Temporal Statistics
We analyze the distribution of terms in GIST based on the publication years of the papers. Table 6 presents the top 10 most frequent terms in awarded papers for each year. This analysis offers insights into the temporal evolution of AI terminology.
Year Terms
2000 feature vector, machine translation, probabilistic model, inference, loss function, feature space, BLEU, translation model, model selection, OOV
2001 reinforcement learning, policy, Ablation study, baseline, machine learning, bias, learning rate, state, gradient descent, tensor
2002 machine learning, machine translation, posterior distribution, inference, weight vector, training set, time complexity, tf-idf, mutual information, dot product
2003 softmax, computer vision, machine learning, local minima, gradient descent, neural network, Dataset, generative model, Gaussian noise, time complexity
2004 deep learning, ground truth, neural network, bounding box, 3D object detection, mean square error, machine learning, cost volume, gradient, receptive field
2005 neural network, loss function, machine translation, learning rate, gradient, perplexity, Kernel, validation set, beam search, feature vector
2006 NLP, machine learning, beam search, language model, Question Answering, Dataset, regularization, overfitting, greedy algorithm, neural network
2007 learning rate, probability distribution, machine learning, deep learning, loss, binary classification, covariance matrix, Adam optimizer, corpus, estimator
2008 machine learning, softmax, dropout, language model, computer vision, bias, baseline, neural network, F1 score, NLP
2009 computer vision, language model, softmax, NLP, deep learning, learning rate, clustering, natural language processing, neural network, loss function
2010 machine learning, learning rate, neural network, gradient descent, Euclidean distance, regularization, convex optimization, learning algorithm, Gaussian noise, gradient
2011 learning rate, NLP, deep learning, overfitting, supervised learning, Lemma, inference, eigenvalue, baseline, Cross Entropy Loss
2012 softmax, NLP, neural network, attention, deep learning, dropout, State-of-the-art, word embedding, learning rate, attention mechanism
2013 machine learning, loss function, Dataset, inference, NLP, learning rate, language model, stochastic gradient descent, probability distribution, decoder
2014 language model, NLP, machine learning, computer vision, convolutional layer, learning rate, deep learning, loss function, softmax, sigmoid
2015 NLP, language model, learning rate, feature vector, baseline, loss function, recall, validation set, probabilistic model, Markov chain
2016 learning rate, decoder, classification, loss function, NLP, Dataset, data augmentation, machine learning, State-of-the-art, beam search
2017 unsupervised learning, supervised learning, neural network, reinforcement learning, Dataset, dynamic programming, inference, validation set, NLP, epoch
2018 NLP, neural network, learning rate, machine learning, language model, Dataset, inference, natural language processing, loss function, BERT
2019 machine learning, learning rate, NLP, neural network, softmax, gradient descent, natural language processing, computer vision, node, Dataset
2020 NLP, Dataset, loss function, language model, natural language processing, learning rate, reinforcement learning, BERT, Adam optimizer, baseline
2021 learning rate, NLP, natural language processing, machine learning, language model, loss function, computer vision, BERT, overfitting, decoder
2022 machine learning, neural network, NLP, learning rate, computer vision, Dataset, loss function, deep learning, supervised learning, node
2023 machine learning, NLP, learning rate, computer vision, deep learning, Dataset, softmax, machine translation, natural language processing, neural network
Table 6: Top 10 most frequent terms in awarded papers for each year from 2000 to 2023.
Conclusion.
The statistical analysis highlights the diversity and interdisciplinary nature of the dataset. Figures 9 and 10 illustrate domain-wise distributions and semantic clusters, while Table 4 quantifies lexical variations across languages. Together, these findings provide a robust understanding of the dataset’s structure, supporting its utility for multilingual and domain-specific AI applications.
B.2 Terminology Collection Details
We selected terms from awarded papers as this approach provides an efficient and manageable way to curate a representative sample of influential AI research. While we acknowledge that valuable AI terminology also exists in non-awarded yet influential papers, capturing all relevant terms across the vast AI field is infeasible. Nonetheless, we believe our dataset offers a comprehensive representation in terms of domain coverage and unique terminology. We considered expanding our selection to include highly cited papers; however, to our knowledge, no automated method reliably identifies such papers. Even if one existed, determining an appropriate citation threshold for inclusion would remain a challenge.
To refine our terminology selection, we employ a two-step procedure to filter out non-nominal phrases. First, we prompt GPT-4 to retain only nouns or noun phrases as candidates. Subsequently, we use a POS tagger to further remove any phrases that do not contain a noun.
We integrate terms from the WikiPedia Glossary of AI777en.wikipedia.org/wiki/Glossary_of_artificial_intelligence, which serves as a comprehensive starting point for artificial intelligence (AI)-related terminology, ensuring alignment with globally recognized concepts and definitions. To enhance multilingual coverage and domain relevance, we also include terms from several other specialized AI terminology initiatives across different languages:
Arabic AI Dictionary888ai.gov.ae/ar/ai-dictionary/: Published by the Arabic Government AI Office, this resource aims to elevate the status of the Arabic language in AI, standardize terminology, reduce linguistic ambiguity, and foster better integration of Arabic speakers into the global AI community. By providing accurate translations and clear definitions for English AI terms, this dictionary promotes knowledge dissemination and encourages collaboration within the Arabic-speaking AI ecosystem.
GRAND_LEXIQUE_FRAN%C3%87AIS
Russian AI Dictionary [^11] https://www.raai.org/pages/
UGFnZVR5cGU6MTAwMw==
This initiative captures the interdisciplinary nature of AI by including terminology drawn from fields such as logic, psychology, linguistics, and cybernetics. Leveraging contributions from Russian and Soviet experts, this dictionary emphasizes the frequency and relevance of terms within AI-specific contexts, providing a culturally adapted yet globally aligned resource for Russian-speaking researchers.
B.3 Terminology Translation Details
Terminology Dataset Collection
Category Selected Conferences
AI AAAI, IJCAI
CV CVPR, ECCV, ICCV
ML ICLR, ICML, NeurIPS, KDD
NLP ACL, EMNLP, NAACL, EACL
LREC, COLING, CoNLL
Web & IR SIGIR, WWW
Table 7: Top AI conferences included in our dataset collection.
Prompt Design for Translation
Arabic Chinese French Japanese Russian
Three-Model Agreement Ratio 10.11% 42.71% 9.86% 16.60% 17.93%
Two-Model Agreement Ratio 30.52% 36.82% 45.15% 40.44% 38.36%
Translation Agreement Analysis
We evaluated the consistency of translations across three models: Claude 3 Sonnet, GPT-3.5-Turbo, and Google Translate API. Table 8 summarizes the agreement ratios for five target languages: Arabic, Chinese, French, Japanese, and Russian. The results indicate significant variation in agreement ratios across languages, with Chinese achieving the highest three-model agreement (42.71%) and Arabic having the lowest (10.11%). These findings underscore the need for human involvement in AI terminology translation, as automatic translations often fall short due to the inherent limitations of current models.
To further assess translation quality, we also examined which model’s outputs align most closely with human annotations by incorporating GPT-4o. Table 9 presents the results, showing that GPT-4o is the most advanced LLM available at the time of this study. This claim is further supported by relevant literature (Wang et al., 2024; Sato et al., 2024; Zhang et al., 2024; OpenAI, 2024). However, our evaluation reinforces the argument that, despite GPT-4o’s advancements, it remains inadequate for precise terminology translation. This highlights the necessity of human annotations to ensure translation accuracy and reliability.
Language Claude 3 vs. Human Google Translate vs. Human GPT-3.5 vs. Human GPT-4o vs. Human (500 Random Examples)
Chinese 69.26% 62.84% 59.20% 76.80%
Arabic 29.26% 36.45% 23.05% 39.80%
French 57.06% 51.44% 14.17% 58.20%
Japanese 57.59% 49.65% 34.07% 67.80%
Russian 39.72% 42.63% 28.23% 41.20%
Table 9: Comparison of translation accuracy between Claude 3, Google Translate, GPT-3.5, and GPT-4o against human translations across different languages.
Human Translation via Mturk
To further validate the translations, we employed human translator through the MTurk platform. Each terminology was translated by 10 independent annotators, resulting in a set of 10 translation candidates for each term. The platform layout and guidance for helping translators give these translations is shown in Figure 7. We compensate annotators in accordance with MTurk’s payment standards.
B.4 Dataset Quality Assessment Details
To evaluate the quality of translations in our dataset, we conducted two tasks involving human annotators. Annotators were presented with pairs of translations generated by different methods and were tasked with evaluating their relative quality. The analysis of their ratings is summarized in Tables 10, 11, and 12. Below, we describe the findings in detail.
In Task 1, we constructed the evaluation dataset by first randomly sampling 1,000 terms per language that had received majority votes. We then filtered out terms where both translation strategies produced identical results, leaving approximately 200 terms per language for human evaluation. Specifically, this resulted in 323 Arabic terms, 185 Chinese terms, 180 French terms, 206 Japanese terms, and 230 Russian terms.
Annotators compared translations selected by GPT-4o (Method 1) with majority-voted translations generated by human annotators (Method 2). The choices were: A. Both translations are good; B. Method 1’s translation is better; C. Method 2’s translation is better; D. Both translations are bad.
Arabic 739, 45.76% 461, 28.54% 329, 20.37% 72, 4.46%
Chinese 468, 50.59% 266, 28.76% 163, 17.62% 25, 2.70%
French 438, 48.67% 274, 30.44% 170, 18.89% 17, 1.89%
Japanese 587, 56.99% 251, 24.37% 159, 15.44% 27, 2.62%
Russian 626, 54.43% 348, 30.26% 150, 13.04% 14, 1.22%
Arabic 376, 46.42% 238, 29.38% 143, 17.65% 46, 5.68%
Chinese 197, 37.17% 228, 43.02% 85, 16.04% 17, 3.21%
French 152, 39.48% 168, 43.64% 53, 13.77% 10, 2.60%
Japanese 295, 57.28% 162, 31.46% 36, 6.99% 21, 4.08%
Russian 172, 39.09% 198, 45.00% 38, 8.64% 25, 5.68%
In Task 2, the number of overlapping terms between the 60-60 evaluation set and our dataset varies across languages for two main reasons. First, some terminology translations appear in certain languages but not others due to the integration of external data sources, as shown in Table 1. Second, some translations are identical in both our dataset and the 60-60 evaluation set; these were removed prior to human evaluation. As a result, the final analysis includes 162 Arabic terms, 106 Chinese terms, 77 French terms, 103 Japanese terms, and 88 Russian terms.
Annotators compared translations from our dataset (Method 1) with those in the 60-60 initiative evaluation set (Method 2). The distribution of choices is shown in Table 11. In this task, a higher percentage of annotators preferred translations from Method 2 (category B) for most languages, especially in Chinese (43.02%), French (43.64%), and Russian (45.00%). However, Japanese translations from Method 1 had a significantly higher percentage in category A, with 57.28% of annotators agreeing that both translations were good. Instances where both translations were rated as bad (category D) remained low across all languages, ranging from 2.60% to 5.68%.
Arabic Chinese French Japanese Russian
Task 1 0.30 0.22 0.37 0.21 0.39
Task 2 0.22 0.41 0.50 0.39 0.41
Inter-Annotator Agreement
Conclusion
B.5 Dataset Coverage Assessment Details
Domain Distribution.
Semantic Clustering.
Coverage Analysis through Rarefaction.
Figure 9: Terminology distribution of the top-6 AI domains in GIST.
Figure 10: UMap visualization of terms in GIST by domain.
Appendix C Additional Experiment Details
C.1 Terminology Refinement via Prompting Details